OpenStudy (anonymous):

Determine whether the given differential equation is exact. If it is exact, solve it. (If it is not exact, enter NOT.) x dy/dx= 2xe^(x) − y + 6x^(2)

OpenStudy (btaylor):

We want to write this in the form \(M(x,y) + N(x,y) \cdot \frac{dy}{dx} = 0\). Then, we take the derivative of M with respect to x and of N with respect to y.